# Early-stopping for Transformer model training

### v1

@@ 3.4 Recognition criteria for Heavy-Tail based on PL fitting
Proof.
$P(\sup_{x\in\mathbb{R}}|\hat{F}_{n}(x)-F(x)|>\epsilon)\leq 2e^{-2n\epsilon^{2}}.$ (3.6)
Consequently, we obtain
$P(d>\epsilon)\leq 2e^{-2n_{\text{tail}}\epsilon^{2}}.$
In practical applications, the PL exponent $\alpha$ is typically unknown and must be estimated from sample data. Maximum Likelihood Estimation (MLE) is a widely used and robust approach for estimating the parameters of a power-law distribution. The MLE for $\hat{\alpha}$ is given by
$\hat{\alpha}=1+n_{\text{tail}}\left(\sum_{i=1}^{n_{\text{tail}}}\ln\frac{x_{i}}{x_{\min}}\right)^{-1},$
Substituting $\hat{\alpha}$ into the CDF yields the empirical KS distance
@@
Proof.
Consider the KS distance
$=d+\sup_{x\geq x_{min}}|F(x;\alpha,x_{min})-F(x;\hat{\alpha},x_{min})|.$ (3.9)
Log-likelihood function is
[…]
Combining these results, we obtain that
$=$ $\left(\frac{x}{x_{min}}\right)^{-\alpha+1}\left(\left(\frac{x}{x_{min}}\right)^{-(\hat{\alpha}-\alpha)}-1\right),$ (3.10)
$F(x;\alpha,x_{min})-F(x;\hat{\alpha},x_{min})=O_{p}\left(\frac{1}{\sqrt{n_{tail}}}\right).$ (3.11)
Hence, the convergence rate of the distance $\hat{d}$ is
$\hat{d}=O_{p}\left(\frac{1}{\sqrt{n_{\text{tail}}}}\right).$
By Proposition 3.1, under the null hypothesis that the ESD of the en.0.s.a.v matrix follows a Power-Law distribution (i.e., exhibits Heavy-Tailed characteristics), the KS distance $\hat{d}$ between the empirical spectral distribution and the fitted Power-Law distribution converges to 0. By defining the corresponding rejection region for the null hypothesis, we can conclude that if the ESD of the weight matrix deviates from the Power-Law form - i.e., in the absence of Heavy-Tailed characteristics– $\hat{d}$ will no longer tend to 0. This observation leads directly to the following corollary.
@@
In the next section, we determined the value of C through Monte Carlo simulation.
3.5 Critical Constant Calibration
Step 1: Set Simulation Parameters
  • Number of Monte Carlo Iterations (k): We set $k=10,000$ repetitions.
Step 2: Generate Synthetic Data
$x=x_{\min}\cdot(1-u)^{-\frac{1}{\alpha_{0}-1}},$
Step 3: Calculate the Normalized Statistic S
  1. 1.
  2. 2.
  3. 3.
    Calculate the normalized KS statistic S:
Step 4: Select Threshold C
Following the steps outlined above, the results of our simulation are presented in Figure 4.
Figure 4: Monte Carlo simulation for calibrating the threshold constant C. Each subplot displays the distribution of the statistic $S=d_{\text{hat}}\cdot\sqrt{n_{\text{tail}}}$ over 10,000 runs, for different true values of $\alpha$ and tail sizes ($n_{\text{tail}}$). The vertical blue line indicates the proposed threshold.
The simulation results indicate that, across 10,000 Monte Carlo iterations for various values of $\alpha$ and tail data sizes ($n_{\text{tail}}$), the upper bound of the normalized KS statistic remains highly stable and does not exceed 2. Consequently, we select C=2 as the critical threshold.
@@ 3.6 Early Stopping criterion and its Efficacy
In this section, we aim to construct a novel spectral criterion to identify the optimal early-stopping epoch for guiding Transformer model training. We find that the Power-Law (PL) fitting-based early-stopping criterion aligns well with the phase-partitioning-based early-stopping criteria described in Section 3.3.
@@
According to Corollary 3.1, the Heavy-Tailed (HT) indicator, defined as the difference $d^{*}-\tilde{d}$ between the threshold distance and the empirical KS distance, can effectively monitor the dynamic evolution of model training and track the formation of Heavy-Tailed spectral structures. The transition of this indicator from negative to positive values signals the entry of the spectral distribution into the Heavy-Tailed regime.
@@
We propose an early-stopping criterion based on $\mathbf{\max_{epoch}\{d^{*}-\tilde{d}\}}$, where the Heavy-Tailed (HT) indicator $d^{*}-d$ reaches its maximum value. This epoch is considered the optimal early-stopping point, at which the model achieves the highest degree of implicit self-regularization and the strongest generalization capability. At this point, the model’s objectives of performance optimization and training time minimization are optimally balanced.
The evolutionary trajectories of Heavy-Tailed indicator across three experimental configurations are illustrated in the figure 5.
@@
Figure 5: Evolutionary Trajectory of the Early-Stopping Metric. (a), (b), and (c) describe the variation of d and d* with training epochs for the three models, respectively; while (b), (d), and (f) illustrate the changes in the difference between d* and d across training epochs for these three models, respectively.
@@
Observation of the evolutionary trajectory of the HT indicator $d^{*}-d$ in Figure 5 shows that, after an initial period of intense adjustment, the indicator rises rapidly, signaling the model’s entry into the Heavy-Tailed phase. Subsequently, the indicator reaches its peak and may exhibit minor fluctuations, suggesting that the Heavy-Tailed characteristics, after attaining their maximum intensity, can slightly diminish during the later stages of training.
Applying the proposed early-stopping criterion, the optimal early-stopping epochs for the three experimental configurations are identified as 100, 100, and 50, respectively. Table 4 presents a comparison of model performance at these early-stopping points against the performance achieved at the final training epoch on the validation set.
Table 4: Validation of Early-Stopping Strategy Effectiveness Based on Spectral Distance Metric: Performance Comparison
@@
Model Early-Stopping Point Final Epoch (205) Relative Difference
Loss Accuracy Loss Accuracy Diff.loss Diff.acc
T1_En-Ch9.60M 2.3812 0.5419 2.3465 0.5512 1.479% -1.687%
T2_En-Ch9.62M 2.4360 0.5322 2.4033 0.5402 1.361% -1.481%
T3_En-Ch19.20M 2.1039 0.5713 2.0248 0.5861 3.907% -1.48%
@@
The results presented in Table 4 clearly demonstrate the effectiveness of the proposed early-stopping strategy. For experiments T1_En-Ch9.60M and T2_En-Ch9.62M, an additional 105 epochs of training yields a loss reduction of less than 1.5% and an accuracy improvement of less than 1.7%. In experiment T3_En-Ch19.20M, extending training by 155 epochs results in a loss reduction of less than 4% and an accuracy gain of less than 1.5%. These findings indicate that the spectral distance-based early-stopping criterion can substantially reduce training time and computational cost while preserving near-final model performance with negligible differences.
@@ 4 Conclusion
The spectral distribution of the self-attention matrix V in the initial encoder layers of Transformers is profoundly influenced by training dynamics and exhibits persistent Heavy-Tailed characteristics during the prolonged phase of stable evolution. We investigate these phenomena from multiple perspectives. Using Random Matrix Theory, we analyze the Power-Law (PL) fitting parameters and their dynamic evolution, revealing that changes in these parameters effectively reflect the progress of model training. Building upon the Heavy-Tailed Self-Regularization theory proposed by Martin and Mahoney (2021) and in contrast to existing spectral early-stopping methods (Meng and Yao, 2023), we partition the training process into three phases and demonstrate that the strong Heavy-Tailed phase (Phase II) constitutes a critical period for effective fine-tuning in Transformer models. This finding offers a novel perspective for understanding the internal training mechanisms of Transformers. Based on these insights, we propose a quantitative criterion for recognizing Heavy-Tailed behavior via PL fitting, which can track the formation dynamics of Heavy-Tailed structures and provide a spectral-based early-stopping criterion for Transformer models. Notably, this approach enables stopping decisions without reliance on a separate validation set.
These findings also raise several compelling questions for future exploration. For example, as Transformer architectures continue to evolve, an important direction is to evaluate whether the proposed spectral early-stopping criterion generalizes to other large language models, such as LLaMA-3 and DeepSeek-MoE. Systematic assessment of the method’s generalizability and limitations represents a crucial avenue for subsequent research. Investigating these underlying mechanisms and extending the criterion will provide significant directions for future work.
@@ Appendix A Theoretical Analysis of the Optimization Dynamics of the Value Matrix $W^{V}$
Recent studies (Gurbuzbalaban et al., 2021) have demonstrated that Heavy-Tailed behavior in deep neural networks is intimately linked to the optimization dynamics induced by gradient descent. In particular, Song et al. (2024) proved that, in a single-layer Transformer, optimizing only the value matrix ($W^{V}$) is sufficient to achieve the global minimum. This finding underscores the dominant and simplifying role of $W^{V}$ during model training.
In practical Transformer training, however, all three attention matrices ($W^{Q}$, $W^{K}$, $W^{V}$) are jointly updated. Due to the non-convexity introduced by the softmax operation, the optimization landscapes associated with $W^{Q}$ and $W^{K}$ are considerably more intricate than that of $W^{V}$. To support our empirical observation that the dynamics of $W^{V}$ exhibit smooth, nearly linear convergence alongside Heavy-Tailed statistical behavior, we extend existing convergence analyses-originally established for quadratic losses-to the cross-entropy loss employed in our translation tasks. Specifically, we analyze the optimization dynamics of $W^{V}$ while keeping $W^{Q}$ and $W^{K}$ fixed.
Our objective is not to establish global convergence for the entire Transformer model but rather to demonstrate that, under mild assumptions, optimizing $W^{V}$ alone follows provably stable gradient dynamics with sublinear convergence rates.
@@ A.3 Smoothness Proof
$\|\nabla_{W^{V}}f(W_{1}^{V})-\nabla_{W^{V}}f(W_{2}^{V})\|_{F}$ $=\|\mathcal{A}^{*}(\nabla_{z}\mathcal{L}_{CE}(z_{1})-\nabla_{z}\mathcal{L}_{CE}(z_{2}))\|_{F}$ (A.6)
$\leq\|\mathcal{A}^{*}\|_{\mathrm{op}}\tfrac{1}{4}\|\mathcal{A}\|_{\mathrm{op}}\|W_{1}^{V}-W_{2}^{V}\|_{F}.$ (A.7)
@@ A.5 Discussion
This analysis demonstrates that when $W^{Q}$ and $W^{K}$ are approximately fixed, the optimization of $W^{V}$ reduces to a convex and $L$-smooth problem, ensuring an $O(1/t)$ convergence rate. This explains the empirically observed stability of $W^{V}$ dynamics and reinforces our conclusion that the Heavy-Tailed phenomenon manifests most prominently in the value matrix.
@@ Appendix B Pervasive Heavy-Tailed Features and Analysis of Representative Matrices
As noted by Clauset et al. (2009) and Alstott et al. (2014), even when a dataset appears to follow a power-law distribution, alternative distributions such as the exponential or log-normal may provide an equally good or even superior fit. To rigorously test this, one can compare the goodness-of-fit among competing models and compute the p-value associated with the power-law hypothesis (Clauset et al., 2009; Klaus et al., 2011; Alstott et al., 2014).
We compared the fits of the Power-Law, Log-Normal, and Exponential distributions to the empirical spectral distributions (ESDs) of the Q, K, and V matrices. In the English–Chinese dataset, we frequently observed that the Log-Normal distribution provided a better fit than the Power-Law—for example, fore.g., for “de.0.a.v”, “de.1.a.v ”, “de.0.s.a.q”, “de.0.s.a.v”, “en.1.s.a.k”, “en.0.s.a.v [^3] Here, ‘de’ denotes the decoder layer, ‘en’ denotes the encoder layer, ‘0’,‘1’,‘2’ indicate the layer indices, ‘a’ represents attention, ‘s.a’ refers to the self-attention layer, and ‘q’, ‘k’, and ‘v’ correspond to the Q, K, V matrix. at epoch=1. Alstott et al. (2014) emphasized that when a plausible generative mechanism exists for the log-normal distribution, it often fits empirical data as well as–or better than–the power-law. Similarly, Clauset et al. (2009) observed that, unless the dataset is exceptionally large, no statistical test can reliably distinguish between the log-normal and power-law fits. Nonetheless, prior work (Martin et al., 2021) has shown that for most weight matrices, both the bulk and the tail of their ESDs can be well described by a power-law distribution. Therefore, we consider the power-law model a reasonable and theoretically grounded choice.
Despite this, we conducted direct comparisons between the power-law and log-normal distributions and obtained results consistent with those of Clauset et al. (2009) and Alstott et al. (2014). Detailed results are omitted due to space constraints. Because the exponential distribution serves as a minimal benchmark for identifying heavy-tailed behavior (Alstott et al., 2014), we further compared the power-law and exponential fits under various parameter settings using the same datasets. By examining the p-values and log-likelihood ratio R, we identified matrices for which the power-law fit significantly outperformed the exponential fit, indicating clear heavy-tailed characteristics. A summary of these results is provided in Table 5.
@@
The results in Table 5 reveal a particularly consistent pattern: the en.0.s.a.v weight matrix persistently exhibits a superior Power-Law (PL) fit compared to the exponential fit, regardless of the dataset or architectural configuration. To further investigate this robust heavy-tailed behavior, we track the evolution of the p-value and the log-likelihood ratio (R) for this matrix across training epochs. The temporal dynamics under different experimental settings are visualized in the figure below.
@@
As evidenced by the p-values and log-likelihood ratios (R), the en.0.s.a.v matrix consistently exhibits pronounced heavy-tailed characteristics throughout training.

### v2

@@ 3.4 Construction of a Power-Law-Based Heavy-Tailed Metric: Key Theoretical Details
The proof of the Lemma is given in Appendix D.1. In practical applications, the PL exponent $\alpha$ is typically unknown and must be estimated from sample data, denoted as $\hat{\alpha}$. The CDF yields the empirical KS distance is
@@
The proof of the proposition is given in Appendix D.2. By Proposition 3.1, under the null hypothesis that the ESD of the en.0.s.a.v matrix follows a Power-Law distribution (i.e., exhibits Heavy-Tailed characteristics), the KS distance $\hat{d}$ between the empirical spectral distribution and the fitted Power-Law distribution converges to 0. By defining the corresponding rejection region for the null hypothesis, we can conclude that if the ESD of the weight matrix deviates from the Power-Law form - i.e., in the absence of Heavy-Tailed characteristics - $\hat{d}$ will no longer tend to 0. This observation leads directly to the following corollary.
@@
We determined the value of C = 2 via Monte Carlo simulation, as detailed in the Appendix E.
@@ 3.5 Early Stopping criterion and its Efficacy
In this section, we aim to construct a novel spectral criterion to identify the optimal early-stopping epoch for guiding Transformer model training. We find that the PL fitting-based early-stopping criterion aligns well with the phase-partitioning-based early-stopping window described in Section 3.3.
@@
According to Corollary 3.1, with the tail threshold fixed at $x_{\min}=0.1$ for our experiments, the Heavy-Tailed (HT) indicator, defined as the difference $d^{*}-\tilde{d}$ between the threshold distance and the empirical KS distance, can effectively monitor the dynamic evolution of model training and track the formation of Heavy-Tailed spectral structures. The transition of this indicator from negative to positive values signals the entry of the spectral distribution into the heavy-tailed regime.
@@
We propose an early-stopping criterion based on $\mathbf{\max_{epoch}\{d^{*}-\tilde{d}\}}$, where the HT indicator $d^{*}-d$ reaches its maximum value. This epoch is considered the optimal early-stopping point, at which the model achieves the highest degree of implicit self-regularization and the strongest generalization capability. At this point, the model’s objectives of performance optimization and training time minimization are optimally balanced.
@@
Figure 4: Evolutionary Trajectory of the Early-Stopping Metric for T1_En-Ch 9.60M. (a) describes the variation of d and d* with training epochs for the three models, respectively; while (b) illustrates the changes in the difference between d* and d across training epochs for these three models, respectively.
@@
The evolutionary trajectories of HT indicator across the experimental configuration T1_En-Ch 9.60M are illustrated in the figure 4, the corresponding plots for T2_En-Ch 9.62M and T3_En-Ch 19.20M are provided in Appendix F. Observation of the evolutionary trajectory of the HT indicator $d^{*}-d$ in Figure 4 shows that, after an initial period of intense adjustment, the indicator rises rapidly, signaling the model’s entry into the Heavy-Tailed phase. Subsequently, the indicator reaches its peak and may exhibit minor fluctuations, suggesting that the Heavy-Tailed characteristics, after attaining their maximum intensity, can slightly diminish during the later stages of training.
@@
Table 4: Early-Stopping Loss, Final Loss (Epoch 205) & Relative Difference Across Models
Model ES Loss Final Loss Re.Diff Loss
T1 2.3812 2.3465 1.479%
T2 2.4360 2.4033 1.361%
T3 2.1039 2.0248 3.907%
@@
Applying the proposed early-stopping criterion identifies the optimal epochs for the three experimental configurations as 100, 100, and 50, respectively. As detailed in Table 4, which compares validation performance at these check-points against the final epoch, the strategy proves highly effective: for experiments T1_En-Ch9.60M and T2_En-Ch9.62M, continuing training for an additional 105 epochs yields a marginal loss reduction of less than 1.5%, while for T3_En-Ch19.20M, extending training by 155 epochs results in less than a 4% improvement. These findings demonstrate that the spectral distance-based criterion can substantially reduce training time and computational costs while preserving near-final model performance with negligible degradation.
@@ 4 Conclusion
The spectral distribution of en.0.s.a.v of Transformer models is profoundly influenced by training dynamics and exhibits persistent Heavy-Tailed (HT) characteristics during the prolonged phase of stable evolution. We investigate these phenomena from multiple perspectives. We analyze the Power-Law (PL) fitting parameters and their dynamic evolution, revealing that changes in these parameters effectively reflect the progress of model training. We partition the training process into three phases and demonstrate that the strong HT phase (Phase II) constitutes a critical period for effective fine-tuning in Transformer models. This finding offers a novel perspective for understanding the internal training mechanisms of LLMs. We propose a quantitative criterion for recognizing HT behavior via PL fitting, which can track the formation dynamics of HT structures and provide a spectral-based early-stopping criterion for Transformer models. Notably, this approach enables stopping decisions without reliance on a separate validation set.
These findings also raise several compelling questions for future exploration. For example, an important direction is to evaluate whether the proposed spectral early-stopping criterion generalizes to other LLMs such as LLaMA-3 and DeepSeek-MoE. Systematic assessment of the method’s generalizability and limitations represents a crucial avenue for subsequent research. Investigating these underlying mechanisms and extending the criterion will provide significant directions for future work.
@@ Appendix A Theoretical Analysis of the Optimization Dynamics of the Value Matrix $W^{V}$
Recent studies (gurbuzbalaban2021heavy) have demonstrated that heavy-tailed behavior in deep neural networks is closely linked to the optimization dynamics induced by gradient descent. In particular, song2024unraveling proved that in a single-layer Transformer, optimizing only the value matrix ($W^{V}$) is sufficient to achieve the global minimum. This finding underscores the dominant and simplifying role of $W^{V}$ during model training.In practical Transformer training, however, all three attention matrices ($W^{Q},W^{K},W^{V}$) are updated jointly. Due to the non-convexity introduced by the softmax operation, the optimization landscapes associated with $W^{Q}$ and $W^{K}$ are considerably more intricate than that of $W^{V}$. To support our empirical observation that the dynamics of $W^{V}$ exhibit smooth, nearly linear convergence alongside heavy-tailed statistical behavior, we extend existing convergence analyses—originally established for quadratic losses—to the cross-entropy loss employed in our tasks. Specifically, we analyze the optimization dynamics of $W^{V}$ while holding $W^{Q}$ and $W^{K}$ fixed.
Our objective is not to establish global convergence for the entire Transformer model, but rather to demonstrate that, under mild assumptions, optimizing $W^{V}$ alone follows provably stable gradient dynamics with sublinear convergence rates.We formalize this theoretical guarantee in the following proposition:
@@
In the following subsections, we provide a detailed analysis to support Proposition LABEL:prop:A1. We begin by formalizing the problem setup in Section A.1. Subsequently, we verify the three claims of the proposition sequentially: Section A.2 establishes the convexity, Section A.3 proves the smoothness condition, and Section A.4 derives the global convergence rate.
@@ A.3 Smoothness Proof
$\|\nabla_{W^{V}}f(W_{1}^{V})-\nabla_{W^{V}}f(W_{2}^{V})\|_{F}$ $=\|\mathcal{A}^{*}(\nabla_{z}\mathcal{L}_{CE}(z_{1})-\nabla_{z}\mathcal{L}_{CE}(z_{2}))\|_{F}$
$\leq\|\mathcal{A}^{*}\|_{\mathrm{op}}\tfrac{1}{4}\|\mathcal{A}\|_{\mathrm{op}}\|W_{1}^{V}-W_{2}^{V}\|_{F}.$
@@ A.5 Discussion
This analysis demonstrates that when $W^{Q}$ and $W^{K}$ are approximately fixed, the optimization of $W^{V}$ reduces to a convex and $L$-smooth problem, ensuring an $\mathcal{O}(1/t)$ convergence rate. This explains the empirically observed stability of $W^{V}$ dynamics and reinforces our conclusion that the heavy-tailed phenomenon manifests most prominently in the $V$ matrix.
@@ Appendix B Pervasive Heavy-Tailed Features and Analysis of Representative Matrices
B.1 Analysis of Representative Matrices
In this section, we rigorously verify the heavy-tailed characteristics of Transformer weight matrices by comparing the goodness-of-fit of the Power-Law distribution against alternative hypotheses. As noted by clauset2009power and alstott2014powerlaw, even when a dataset appears to follow a power-law distribution, alternative distributions such as the exponential or log-normal may provide an equally good or even superior fit. To rigorously test this, one can compare the goodness-of-fit among competing models and compute the p-value associated with the power-law hypothesis (clauset2009power; klaus2011statistical; alstott2014powerlaw).
We compared the fits of the Power-Law, Log-Normal, and Exponential distributions to the empirical spectral distributions (ESDs) of the Q, K, and V matrices. In the English–Chinese dataset, we frequently observed that the Log-Normal distribution provided a better fit than the Power-Law—for example, fore.g., for “de.0.a.v”, “de.1.a.v ”, “de.0.s.a.q”, “de.0.s.a.v”, “en.1.s.a.k”, “en.0.s.a.v [^4] Here, ‘de’ denotes the decoder layer, ‘en’ denotes the encoder layer, ‘0’,‘1’,‘2’ indicate the layer indices, ‘a’ represents attention, ‘s.a’ refers to the self-attention layer, and ‘q’, ‘k’, and ‘v’ correspond to the Q, K, V matrix. at epoch=1. alstott2014powerlaw emphasized that when a plausible generative mechanism exists for the log-normal distribution, it often fits empirical data as well as–or better than–the power-law. Similarly, clauset2009power observed that, unless the dataset is exceptionally large, no statistical test can reliably distinguish between the log-normal and power-law fits. Nonetheless, prior work (martin2021predicting) has shown that for most weight matrices, both the bulk and the tail of their ESDs can be well described by a power-law distribution. Therefore, we consider the power-law model a reasonable and theoretically grounded choice.
Despite this, we conducted direct comparisons between the power-law and log-normal distributions and obtained results consistent with those of clauset2009power and alstott2014powerlaw. Detailed results are omitted due to space constraints. Because the exponential distribution serves as a minimal benchmark for identifying heavy-tailed behavior (alstott2014powerlaw), we further compared the power-law and exponential fits under various parameter settings using the same datasets. By examining the p-values and log-likelihood ratio R, we identified matrices for which the power-law fit significantly outperformed the exponential fit, indicating clear heavy-tailed characteristics. A summary of these results is provided in Table 5.
@@ B.1 Analysis of Representative Matrices
The results in Table 5 reveal a particularly consistent pattern: the en.0.s.a.v weight matrix persistently exhibits a superior Power-Law (PL) fit compared to the Exponential fit, regardless of the dataset or architectural configuration. To further investigate this robust heavy-tailed behavior, we tracked the evolution of the p-value and the log-likelihood ratio ($R$) for this matrix across training epochs. The temporal dynamics under different experimental settings are visualized in the figures below. Specifically, we identify the condition where the p-value is <0.1<0.1 and the log-likelihood ratio is positive ($R>0$) as indicative of pronounced heavy-tailed characteristics. Based on this criterion, the trends observed in Figures 5, 6 lead to the final conclusion that the en.0.s.a.v matrix consistently exhibits significant heavy-tailed characteristics throughout the training process.
@@
B.2 Supplementary experiments with three training phases
Figure 7: Evolution of $\alpha$ and $x_{\text{min}}$ for en.0.s.a.v matrix tracked across T1_En-Ch9.60M, T2_En-Ch9.62M.
Figure 7 illustrates the detailed trajectories of the fitted parameters ($\alpha$ and $x_{\min}$) for the en.0.s.a.v matrix as a function of training epochs, specifically under the experimental configurations T1_En-Ch9.6M and T2_En-Ch9.6M.
To further substantiate the correspondence between these internal spectral dynamics and external learning performance, Figure 8 depicts the evolution of Loss and Accuracy across all training epochs for the three experimental setups. Given that classic early stopping is typically implemented by monitoring validation set performance, this comparison is essential. During Phase I, the model exhibits the most rapid performance improvement, characterized by a sharp decline in loss and a swift increase in accuracy. In Phase II, the rate of improvement decelerates, and the curves smoothen as the model undergoes steady learning and refinement. Finally, in Phase III, the performance curves plateau, with both validation loss and accuracy showing negligible further gains. This saturation aligns consistently with the microscopic phenomena observed in the spectral tail.